OpenStudy (anonymous):

Simplify the factorial Expression: (2n-1)!/(2n+1)! How do you solve it??

Nnesha (nnesha):

subtract one until you get same parenthesesat the denominator & numerator to cancel out :D

OpenStudy (anonymous):

can you show me please

Nnesha (nnesha):

\[\huge\rm \frac{ (2n-1)! }{ (2n+1)! }\] (2n-1)-1 or (2n +1) -1

Nnesha (nnesha):

keep subtracting one until you can cancel something :) is that make sense ? seems like nope :

OpenStudy (anonymous):

where are you getting the one from

Nnesha (nnesha):

that's the rule you have to subtract 1 for example if there are numbers \[\huge\rm \frac{ 4! }{ 3! } = \frac{ 4 \times 3 \times 2 \times 1 }{ 3 \times 2 \times 1 }\] 4-1 =3 3-1=2 2-1=1

OpenStudy (anonymous):

so (2(-1)-1)-1/(2(-1)+1)-1

Nnesha (nnesha):

nope you don't have to replace n by -1 \[\large\rm (2n+1)\color{red}{-1} = 2n +1\color{red}{-1}\]

Nnesha (nnesha):

combine like terms +1-1 = 0 so 2n left \[\rm \frac{ (2n-1)! }{ (2n+1)2n! }\] now subtract again (2n ) -1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/2n(2n+1)

Nnesha (nnesha):

how did you get 1 at the numerator ?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

thats the answer in the back of the book

Nnesha (nnesha):

nope like i said you have to cancel something so to do that subtract from (2n+1)!

Nnesha (nnesha):

yep but i'm asking how did you get :)

OpenStudy (anonymous):

i dont understand :(

Nnesha (nnesha):

\[\rm \frac{ (2n-1)! }{ (2n+1\color{red}{-1})(2n\color{red}{-1})! }\] subtract one

Nnesha (nnesha):

ik :( wait... i should tag some factorial experts...

OpenStudy (anonymous):

ok

Nnesha (nnesha):

@amistre64

OpenStudy (anonymous):

dont worry about it ill just ask the teacher about it

Nnesha (nnesha):

 \[(2n+1\color{red}{-1}) = 2n\] then subtract one again \[(2n)\color{Red}{-1} = (2n-1)\] now 2n-1 is at the denominator and at the numerator so you cancel out that